OpenStudy (anonymous):

the sum of Jane and Carlas ages is 20. in 5 years time jane will be twice as old as Carla. How old are they now?

OpenStudy (anonymous):

jane is 15 and carla is 5

OpenStudy (dumbcow):

wait i might be wrong

OpenStudy (anonymous):

j+c=20 j+5=2(c+5) 20-c+5=2c+10 3c=15 c=5 j=15

OpenStudy (anonymous):

I laughed at the age with decimals. LOL

OpenStudy (anonymous):